# 3D-LATTE: Latent Space 3D Editing from Textual Instructions

### v1

@@ 3 Methodology
where $CT$ is an alignment function that maps a token index from $p^{\ast}$ to the corresponding token index in $p$. In the case where there is no match, it returns $None$. By also injecting self-attention during the early timesteps of the denoising process, we encourage the model to preserve 3D coarse structural relationships, such as part layout and spatial symmetry, before the introduction of semantic details. In the case of self-attention, the $Edit$ function is defined as follows:
$\hat{A}_{t}^{\text{self}}=Edit(A^{\text{self}},(A^{\ast})^{\text{self}},t):=\begin{cases}(A_{t}^{\ast})^{\text{self}}&\text{if }t<\tau_{\text{self}}\\ A_{t}^{\text{self}}&\text{otherwise}\end{cases}$
@@
where $\tau_{\text{self}}$ denotes the timestep until which injection is applied.
@@ 3.4 Mask Generation and Region Editing
To extract the target edit area, we formulate a query to a large vision-language model (GPT-4o) that includes the original prompt, the edit instruction, and a rendered front-view of the object. The model is asked to identify which part of the object is affected by the edit (e.g., “the shirt of the teddy bear”). We use these identified editing regions to generate 2D multi-view consistent segmentation masks. More specifically, GroundingDINO Liu et al. (2024b) processes the multi-view rendered images and the edit regions, proposed by GPT-4o to generate bounding boxes, which are subsequently refined into 2D segmentation masks tracked by SAM 2 Ravi et al. (2024). This process is outlined in the supplementary. Benefiting from our pixel-aligned representation, these masks naturally define a 3D segmentation of the corresponding pixel-aligned Gaussians. To allow for geometric modifications and greater flexibility, we approximate the initial 3D mask expansion by computing the averaged cross-attention map $A_{t,w^{\ast}}$ for the prompt token $w^{\ast}$ that refers to the desired editing region (e.g., “tutu” in Figure 2), and applying a threshold to obtain binary masks that indicate 3D regions influenced by the new concept. The final 3D editing region $M$ is defined as the union of the original mask predicted by SAM 2 lifted to 3D and the thresholded attention-derived mask. We constrain the diffusion process to this area, by blending the original latent $z_{t-1}$ with the edited latent $z_{t-1}^{*}$. Formally,
@@
where $\odot$ denotes element-wise multiplication. Our method also seamlessly supports user-defined masks $M_{user}$, to allow for greater controllability.
@@ 3.5 Geometry Regularization
The editing signal via attention injection can introduce some uncertainty in the editing regions, resulting in Gaussians prone to semi-transparencies and premature collapse. We find that a simple geometric regularizer can help circumvent these artifacts. Thus, we incorporate a soft geometry-aware classifier guidance mechanism that stabilizes the optimization by softly penalizing the removal of Gaussians in regions that are highly relevant to the edit. To this end, we compute a soft mask $R_{t}^{i}\in[0,1]$ for each Gaussian $i$, which reflects how relevant that Gaussian is to the current edit at denoising step $t$. We define the editing signal at each timestep as the L1 difference between the noise predictions conditioned on the edited and original prompts:
$D^{i}=\left\|\epsilon_{\theta}(z_{t},t,p^{\ast})-\epsilon_{\theta}(z_{t},t,p)\right\|_{1},$
where $\epsilon_{\theta}(z_{t},t,p)$ and $\epsilon_{\theta}(z_{t},t,p^{\ast})$ denote the denoising model’s noise predictions under the original and edited prompts, respectively. To obtain a soft mask $R_{t}$ that localizes the highly relevant Gaussians in each denoising step, we normalize the absolute difference values $D^{i}$ across all Gaussians via global min–max normalization. Thus, $R_{t}^{i}$ denotes the relevance weight of the $i$-th Gaussian, arranged in a pixel-aligned grid of size $V\times H\times W$. Our intuition is that the magnitude of this mismatch reflects the relevance of each 3D Gaussian to the edit.
As the existence of Gaussians is determined by their opacity and covariance, our regularization loss is defined as:
$\mathcal{L}_{\text{geo}}=\lambda_{o}\sum_{i}R_{t}^{i}\cdot\exp(-\gamma_{o}\cdot o_{i})+\lambda_{\Sigma}\sum_{i}R_{t}^{i}\cdot\exp(-\gamma_{\Sigma}\cdot\operatorname{Tr}(\Sigma_{i}))$ (6)
where $o$ and $\Sigma$ are obtained by decoding the denoised gaussian splat latents with the VAE decoder $D_{\phi}$ from Lin et al. (2025). Here, $\gamma_{o}$ and $\gamma_{\Sigma}$ control the sharpness of the penalties, and $\lambda_{o}$, $\lambda_{\Sigma}$ weigh the opacity and scale terms. This formulation softly penalizes low opacity and insufficient spatial support. We incorporate our geometry-aware regularization loss as a guidance signal during denoising:
@@ 3.6 Frequency Annealing
The injection of attention maps from the source prompt during editing can impact the model’s denoising ability, intensifying the misalignment between the condition and diffusion spaces. In some examples, this causes the model to overemphasize high-frequency visual features of the source, such as logos or decorative textures. These features may become partially preserved and degrade into artifacts on the surface of the generated 3D asset. We provide illustrations in the ablation study.
Consistent with similar observations in 2D Si et al. (2024), in 3D generation, high-frequency components correspond to fine details while low-frequency components primarily capture global geometric structure and semantic layout. Thus, we introduce a frequency annealing strategy, where we apply spectral modulation in the Fourier domain of the skip feature maps from the U-Net skip connections at each denoising step. Given the skip feature map $h_{l,t}\in\mathbb{R}^{V\times C\times H\times W}$ at layer $l$ and timestep $t$, we apply the following steps:
$\begin{array}[]{ll}\begin{aligned} F(h_{l,t})&=\text{FFT}(h_{l,t})\\ F^{\prime}(h_{l,t})&=F(h_{l,t})\odot\beta_{l,t}\\ h^{\prime}_{l,t}&=\text{IFFT}(F^{\prime}(h_{l,t}))\end{aligned}&\quad\begin{aligned} \beta_{l,t}(r)=\begin{cases}s_{l},&\text{if }t>\tau\text{ and }r<r_{\text{thresh}}\\ s_{h},&\text{if }t\leq\tau\text{ and }r\geq r_{\text{thresh}}\\ 1,&\text{otherwise}\end{cases}\end{aligned}\end{array}$ (8)
Here, $\text{FFT}(\cdot)$ and $\text{IFFT}(\cdot)$ denote the Fourier transform and its inverse. The modulation mask $\beta_{l,t}$ applies scaling based on the radius $r$. This encourages the model to boost low-frequency components during early denoising with attention injection and then transition to emphasize high-frequency components, enabling structural preservation of the source without unwanted oversmoothed textures.
@@
Many existing 3D and multi-view generation models, particularly those fine-tuned on datasets with plain surface details, are constrained to low-resolution outputs and struggle to capture fine-grained geometric and appearance details. In contrast, our target is high-fidelity asset generation. However, in challenging cases involving finer structures when rendering our learned 3D Gaussian representation at higher resolutions, we observe a degradation in appearance quality, characterized by blurred textures and limited details. Drawing inspiration from Haque et al. (2023) to improve visual fidelity while preserving 3D structure and consistency, we leverage an iterative dataset update technique. Unlike prior work that primarily uses such pipelines for editing Haque et al. (2023); Chen et al. (2023), we adapt them for 3D enhancement. Specifically, we follow an iterative process that comprises three key steps: i) rendering high-resolution views from the edited 3DGS representation, ii) enhancing these views by inputting noised version of them to the 2D diffusion backbone model, conditioned on the original edited images and , (iii) re-optimizing the 3DGS model with the updated enhanced images. To constrain the extent of correction to the editing area, the updated image is represented as: $I_{\text{blend}}=M\odot I_{e}+(1-M)\odot I_{\text{src}},$ where $M$ denotes the mask containing the changed region, $I_{e}$ denotes the enhanced image and $I_{\text{src}}$ denotes the original image of the source object. This process is outlined in the supplementary material.
As our 2D enhancement backbone, we leverage the ControlNet-Tile Zhang et al. (2023) model, which is designed for structure-preserving super-resolution, and is effective at restoring high-frequency details and resolving visual artifacts. By progressively re-rendering and updating the images we converge to a globally consistent higher-fidelity 3D representation.
@@ 4 Experiments
Table 1: Quantitative comparison using directional CLIP score metrics.
Method CLIP Dir CLIP Diff No-Edit
MVEdit Chen et al. (2024b) 0.11 0.080
Vox-E Sella et al. (2023) 0.12 0.053
GaussCTRL Wu et al. (2024a) 0.07 0.033
Ours 0.18 0.040
To evaluate our method, we construct a benchmark comprising 20 diverse 3D assets, each paired with three distinct text edit instructions, resulting in a total of 60 edited samples. The assets are sourced from the Objaverse Deitke et al. (2023b) and Google Scanned Objects (GSO) Downs et al. (2022) datasets. As baselines, we include Vox-E Sella et al. (2023), a voxel-based method using SDS, MVEdit Chen et al. (2024b) and GaussCTRL Wu et al. (2024a). For the quantitative evaluation, we adopt the CLIP directional similarity (CLIP-dir) metric Haque et al. (2023) to quantify semantic alignment to edits by measuring direction changes in CLIP text embeddings and corresponding image embeddings of multi-view renderings. To measure how well unedited regions are preserved, we include the CLIP-diff-no-edit metric Erkoç et al. (2024), which measures the CLIP score difference between input and output images using a modified text prompt where the edited part is replaced with a generic placeholder. Lower values indicate better shape preservation. More details are presented in the supplementary. To complement these metrics, we report results from GPTEval3D Wu et al. (2024b), a GPT-4V-based evaluation protocol in which the model compares multi-view renderings from different methods along three criteria: Text Prompt–3D Alignment, 3D Plausibility, and Texture Details.
@@ 4.1 Qualitative Results
As illustrated in Figure 3 and 5, our method produces high-quality edits that remain faithful to the input prompt. It is capable of handling significant geometric transformations, such as turning a shovel into a flower, or a house into a Greek ancient building while preserving unedited regions and the structural integrity of the original object. As shown in Figure 3, MVEdit Chen et al. (2024b) successfully modifies appearance but struggles with morphological changes. We also observe multi-view inconsistencies, such as two faces on the teddy bear. GaussCTRL Wu et al. (2024a) exhibits similar limitations as it also struggles with geometric transformations, due to its reliance on depth-guidance. Vox-E Sella et al. (2023) handles geometric changes more robustly but suffers from low quality and its mechanism for edit mask detection sometimes fails to localize edits precisely (e.g., wings on the character or shovel). Moreover, it is prone to unnatural 3D geometries due to the limited 3D awareness of SDS. Additional comparisons can be found in the supplementary.
@@ 4.2 Quantitative Results
The quantitative comparison in Table 1 supports the qualitative findings that our method surpasses the baselines in terms of semantic alignment of predictions to the edited text prompts measured by the CLIP-dir score. The CLIP-no-diff score demonstrates that we preserve shapes effectively, compared to the other methods and we achieve a balanced tradeoff between editing and shape preservation. While GaussCTRL achieves a lower Diff-No-Edit score, this can be attributed to its frequent failure to apply the intended edits, reflected in its significantly lower CLIP-dir score. In Table 2, we show the results of the GPTEval3D Wu et al. (2024b) metric in which GPT-4V shows a distinct preference for our method.
       Input MVEdit Chen et al. (2024b) Vox-E Sella et al. (2023) GaussCTRL Wu et al. (2024a) Ours
We also conduct a user study with 57 participants, who were asked to choose the best editing results among ours and the three baselines, based on two criteria: visual quality and faithfulness to the edit prompt. The results in Figure 4 show that our method was preferred by a significant margin.
Table 2: Quantitative Comparison using GPTEval3D Wu et al. (2024b). Following Erkoç et al. (2024) the scores represent the percentage of pairwise comparisons in which our method was favored over the respective baselines.
@@
MVEdit Chen et al. (2024b) 89% 72% 71%
Vox-E Sella et al. (2023) 77% 78% 79%
GaussCTRL Wu et al. (2024a) 96 % 84% 84%
@@ 4.3 Ablation study
Effect of 3D Enhancement.
Figure 6(a) showcases the effect of our proposed 3D enhancement module. The top/bottom rows show a rendered view of the 3D objects before and after enhancement, demonstrating that the enhancement module enhances details and sharpens textures while maintaining structural coherence (e.g., sharper architectural details in the left example).
Effect of Geometry Regularization.
We also qualitatively demonstrate the effectiveness of our geometry regularization guidance term. As shown in Figure 6(b), without this term, some edited regions become partially transparent or vanish entirely, leading to degraded geometry. The proposed regularization mitigates these issues, resulting in more complete edits.
Effect of Frequency Annealing.
As shown in Figure 6(c), without frequency annealing, complex patterns in the source object, such as logos or prints that contain high-frequency information, can be overemphasized by the model, resulting in noisy or corrupted textures in the final edit.
       (a) 3D Enhancement. (b) Geometry Regularization. (c) Frequency Annealing.
Figure 6: Qualitative Ablation Study. We illustrate visual results showing the impact of the core components of our pipeline. We encourage the reader to zoom in.

### v3

@@ 3 Method
By also injecting 3D self-attention during the early timesteps of the denoising process, we encourage the model to further preserve 3D structural relationships of the source asset, such as part layout, composition and spatial symmetry, before the introduction of semantic details. As illustrated in Figure 4, the spectral decomposition of the 3D self-attention graph reflects 3D scene composition and groups 3D Gaussians into distinct semantic parts. In the case of self-attention, the $F_{self}^{3D}$ function is defined as follows:
$\hat{W_{\mathcal{G}}}_{t}^{\text{self}}$ $=F_{\text{self}}^{3D}\!\left(W_{\mathcal{G}}^{\text{self}},(W_{\mathcal{G}}^{\ast})^{\text{self}},t\right)$ (3)
[…]
@@
where $\tau_{\text{self}}$ is the timestep until which injection is applied.
@@ 3.4 Mask Generation and Region Editing
To extract the target edit area, we formulate a query to a large vision-language model (GPT-4o) that includes the original prompt, the edit instruction, and a rendered front-view of the object. The model is asked to identify which part of the object is affected by the edit (e.g., “the shirt of the teddy bear”). We use these identified editing regions to generate 2D multi-view consistent segmentation masks. More specifically, GroundingDINO [liu2023grounding] processes the multi-view rendered images and the edit regions, proposed by GPT-4o to generate bounding boxes, which are subsequently refined into 2D segmentation masks tracked by SAM 2 [ravi2024sam2]. This process is outlined in the supplementary. Benefiting from our pixel-aligned representation, these masks naturally approximate a 3D segmentation of the corresponding pixel-aligned Gaussians. To allow for flexible geometric modifications, we approximate the initial 3D mask expansion by computing the cross-attention map $\overline{{W_{\mathcal{G}}}_{t}^{r^{\ast}}}$ averaged over all time steps for the prompt token $r^{\ast}$ that refers to the desired editing region (e.g., “tutu” in Figure 2), and applying a threshold to obtain binary masks that indicate 3D regions influenced by the new concept. The final 3D editing region $M$ is defined as the union of the original masks predicted by SAM 2 lifted to 3D and the attention-derived mask. We constrain the diffusion process to this area, by blending the original latent $z_{t-1}$ with the edited latent $z_{t-1}^{*}$ as follows:
@@
where $\odot$ denotes element-wise multiplication. Our method also supports user-defined masks.
@@ 3.5 Geometry Regularization
The editing signal via attention injection can introduce some uncertainty in the editing regions, resulting in Gaussians prone to semi-transparencies and premature collapse. We find that a simple geometric regularizer can help circumvent these artifacts. Thus, we incorporate a soft geometry-aware classifier guidance mechanism that softly penalizes the removal of Gaussians in regions that are highly relevant to the edit. To this end, we compute a soft mask $R_{t}^{i}\in[0,1]$ for each Gaussian $i$, which reflects how relevant that Gaussian is to the current edit at denoising step $t$. We define the editing signal at each timestep as the L1 difference between the noise predictions conditioned on the edited and original prompts: $D^{i}=\left\|\epsilon_{\theta}(z_{t},t,p^{\ast})-\epsilon_{\theta}(z_{t},t,p)\right\|_{1},$ where $\epsilon_{\theta}(z_{t},t,p)$ and $\epsilon_{\theta}(z_{t},t,p^{\ast})$ denote the denoising model’s noise predictions under the original and edited prompts, respectively. To obtain a soft relevance mask $R_{t}$ we normalize the absolute difference values $D^{i}$ across all Gaussians via global min–max normalization. Thus, $R_{t}^{i}$ denotes the relevance weight of the $i$-th Gaussian, arranged in a pixel-aligned grid of size $V\times H\times W$. Our intuition is that the magnitude of this mismatch reflects the relevance of each 3D Gaussian to the edit. As the existence of Gaussians is determined by their opacity and covariance, our regularization loss is defined as:
$\mathcal{L}_{\text{geo}}=\lambda_{o}\sum_{i}R_{t}^{i}\cdot\exp(-\gamma_{o}\cdot o_{i})+\lambda_{\Sigma}\sum_{i}R_{t}^{i}\cdot\exp(-\gamma_{\Sigma}\cdot\operatorname{Tr}(\Sigma_{i}))$
where $o$ and $\Sigma$ are obtained by decoding the denoised gaussian splat latents with the VAE decoder $D_{\phi}$ from [lin2025diffsplat]. Here, $\gamma_{o}$ and $\gamma_{\Sigma}$ control the sharpness of the penalties, and $\lambda_{o}$, $\lambda_{\Sigma}$ weigh the opacity and scale terms. This formulation softly penalizes low opacity and insufficient spatial support. We incorporate our geometry-aware regularization loss as a guidance signal during denoising:
@@ 3.6 Frequency Annealing
The injection of attention from the source prompt during editing can impact the model’s denoising ability, intensifying the misalignment between the condition and diffusion spaces. In some examples, this causes the model to overemphasize high-frequency visual features of the source, such as decorative textures. These may become partially preserved and degrade into artifacts on the surface of the generated 3D asset. Consistent with observations in 2D [wu2024freeinit, si2023freeu], in 3D generation, high-frequency components correspond to fine details while low-frequencies primarily capture 3D geometric structure and semantic layout. Thus, we introduce a frequency annealing strategy, where we apply spectral modulation in the Fourier domain of the feature maps from the U-Net skip connections at each denoising step. Given the gaussian feature map $h_{l,t}\in\mathbb{R}^{V\times C\times H\times W}$ at layer $l$ and timestep $t$, we apply the following steps:
$\begin{aligned} &\begin{aligned} F(h_{l,t})&=\mathrm{FFT}(h_{l,t})\\ F^{\prime}(h_{l,t})&=F(h_{l,t})\odot\beta_{l,t}\\ h^{\prime}_{l,t}&=\mathrm{IFFT}(F^{\prime}(h_{l,t}))\end{aligned}&&\quad\begin{aligned} \beta_{l,t}(r)&=\begin{cases}s_{l},&\text{if }t>\tau\text{ and }r<r_{\text{thresh}},\\[2.0pt] s_{h},&\text{if }t\leq\tau\text{ and }r\geq r_{\text{thresh}},\\[2.0pt] 1,&\text{otherwise.}\end{cases}\end{aligned}\end{aligned}$
Here, $\text{FFT}(\cdot)$ and $\text{IFFT}(\cdot)$ denote the Fourier transform and its inverse. The modulation mask $\beta_{l,t}$ applies scaling based on the radius $r$ and the terms $s_{l}$ and $s_{h}$ denote the scaling factors applied to the low- and high-frequency components respectively. This encourages the model to boost low-frequency components during early denoising with attention injection and then transition to emphasize high-frequency components, enabling structural preservation of the source without unwanted oversmoothed textures.
@@
Many existing text-to-3D generation models, particularly those fine-tuned on datasets with plain surface details, are constrained to lower-resolution outputs and struggle to capture fine-grained geometric and appearance details. Thus, in challenging cases involving finer structures when rendering our 3D Gaussian representation at higher resolutions, we observe a degradation in appearance quality, characterized by blurred textures and limited details. Drawing inspiration from [Haque2023instructnerf] to improve visual fidelity while preserving 3D structure and consistency, we leverage an iterative dataset update technique. Unlike prior work that primarily uses such pipelines for editing [Haque2023instructnerf, chen2023gaussianeditor], we adapt them for 3D enhancement. Specifically, we follow an iterative process that comprises three key steps: i) rendering high-resolution views from the edited 3DGS representation, ii) enhancing these views by inputting noised version of them to the 2D diffusion backbone model, conditioned on the original edited images and, (iii) re-optimizing the 3DGS model with the updated enhanced images. To constrain the extent of correction to the editing area, the updated image is represented as: $I_{\text{blend}}=M\odot I_{e}+(1-M)\odot I_{\text{src}},$ where $M$ denotes the mask containing the changed region, $I_{e}$ denotes the enhanced image and $I_{\text{src}}$ denotes the image of the source object. This process is shown in the supplementary.
As our 2D enhancement backbone, we leverage the ControlNet-Tile model [zhang2023adding], which is designed for structure-preserving super-resolution, and is effective at restoring high-frequency details. By progressively re-rendering and updating the images we converge to a globally consistent higher-fidelity 3D representation.
@@ 4 Experiments
Method CLIP Dir CLIP Diff No-Edit CLIP-Dir-Con
[…]
PDS [Koo2024PDS] 0.051 0.094 0.55
InstructGS2GS [igs2gs] 0.069 0.082 0.64
[…]
Table 1: Quantitative comparison using CLIP score metrics. Best, second-best and third-best results are indicated.
To evaluate our method, we construct a benchmark of 25 diverse 3D assets, each paired with a set of distinct edit instructions, resulting in a total of 100 samples. The assets are sourced from the Objaverse [Deitke2023objaverse] and Google Scanned Objects (GSO) [Downs2022GSO] datasets. As main baselines, we include Vox-E [sella2023vox], a voxel-based method using SDS, MVEdit [mvedit2024], GaussCTRL [wu2024gaussctrl] and Edit360 [huang2025edit360]. For completeness, we also report quantitative results on two additional well-studied baselines: InstructGS2GS [igs2gs] and Posterior Distillation Sampling (PDS) [Koo2024PDS]. For the quantitative evaluation, we adopt the CLIP directional similarity (CLIP-dir) metric [Haque2023instructnerf] to quantify semantic alignment to edits by measuring direction changes in CLIP text embeddings and corresponding image embeddings of multi-view renderings. To measure how well unedited regions are preserved, we include the CLIP-diff-no-edit metric [erkoc2024preditor3d], which measures the CLIP score difference between input and output images using a modified text prompt where the edited part is replaced with a generic placeholder. Lower values indicate better shape preservation. Finally, to quantitatively evaluate if the edit propagates consistently across viewpoints we adopt the CLIP direction consistency score (CLIP-Dir-Con) [Haque2023instructnerf], which computes how much the editing directions differ across frames. More details about the metrics are presented in the supplementary. To complement these metrics, we report results from GPTEval3D [wu2023gpteval3d], a GPT-4V-based evaluation protocol in which the model compares multi-view renderings along three criteria: Text Prompt–3D Alignment, 3D Plausibility, and Texture Details.
@@ 4.1 Qualitative Results
As illustrated in Figure 6 and 5, our method produces high-quality, multi-view consistent edits that are faithful to the input prompt. It is capable of handling significant geometric transformations, such as turning a shovel into a flower, or a house into an ancient building while preserving unedited regions and the structural integrity of the original object. As shown in Figure 5, MVEdit [mvedit2024] successfully modifies appearance but struggles with morphological changes and suffers from multi-view inconsistencies. GaussCTRL [wu2024gaussctrl] exhibits similar limitations as it also struggles with geometric transformations, due to its reliance on depth-guidance. Vox-E [sella2023vox] handles geometric changes more robustly but suffers from low quality and sometimes fails to localize edits (e.g., wings on the character). Moreover, it is prone to unnatural 3D geometries due to the limited 3D awareness of SDS. Edit360 [huang2025edit360] shows promise in edit faithfulness, however, the fusion of front and edited view trajectories introduces significant multi-view inconsistencies. Additional comparisons can be found in the supplementary.
@@ 4.2 Quantitative Results
The quantitative comparison in Table 1 supports the qualitative findings that our method surpasses the baselines in terms of semantic alignment of predictions to the edit prompts measured by the CLIP-dir score. The CLIP-no-diff score demonstrates that we preserve shapes effectively, compared to the other methods and we achieve a balanced tradeoff between editing and shape preservation. While GaussCTRL achieves a lower Diff-No-Edit score, this can be attributed to its frequent failure to apply the edits, leaving the objects unaltered. This is reflected in its significantly lower CLIP-dir score and in the qualitative examples. In Table 2, we show the results of the GPTEval3D [wu2023gpteval3d] metric in which GPT-4V shows a distinct preference for our method.
We also conduct a user study with 57 participants, who were asked to choose the best editing results among ours and three baselines, based on two criteria: visual quality and faithfulness to the edit prompt. The results in Figure 7 show that our method was preferred by a significant margin.
@@
MVEdit [mvedit2024] 87% 71% 70%
Vox-E [sella2023vox] 78% 81% 78%
GaussCTRL [wu2024gaussctrl] 94% 83% 81%
Edit360 [huang2025edit360] 67% 90% 72%
Table 2: Quantitative comparison using GPTEval3D. Following [erkoc2024preditor3d], the scores represent the percentage of comparisons where our method wins over the respective baselines.
@@ 4.3 Ablation study
We illustrate visual results showing the impact of the core components of our pipeline. Additional quantitative results are included in the supplementary material.
Effect of 3D Enhancement. Figure 8 showcases the effect of our proposed 3D enhancement module. The top/bottom rows show a rendered view of the 3D objects before and after enhancement, demonstrating that the enhancement module recovers fine details and sharpens textures while maintaining 3D geometry(e.g., sharper architectural details).
Figure 8: Effect of 3D enhancement.
Effect of Geometry Regularization. We also qualitatively demonstrate the effectiveness of our geometry regularization guidance term. As shown in Figure 9(a), without this term, some edited regions become partially transparent or vanish entirely, leading to degraded geometry. Our regularization mitigates these issues, resulting in robust edits.
Effect of Frequency Annealing. As shown in Figure 9(b), without frequency annealing, complex patterns in the source object, such as logos or prints that contain high-frequency information, can be overemphasized by the model, resulting in noisy textures in the final edit.
(a) Geometry Regularization. (b) Frequency Annealing.
Figure 9: Effect of our proposed additional components.